# Geometry-Informed Neural Networks

### v3

@@ Abstract
Geometry is a ubiquitous tool in computer graphics, design, and engineering. However, the lack of large shape datasets limits the application of state-of-the-art supervised learning methods and motivates the exploration of alternative learning strategies. To this end, we introduce geometry-informed neural networks (GINNs) – a framework for training shape-generative neural fields without data by leveraging user-specified design requirements in the form of objectives and constraints. By adding diversity as an explicit constraint, GINNs avoid mode-collapse and can generate multiple diverse solutions, often required in geometry tasks. Experimentally, we apply GINNs to several validation problems and a realistic 3D engineering design problem, showing control over geometrical and topological properties, such as surface smoothness or the number of holes. These results demonstrate the potential of training shape-generative models without data, paving the way for new generative design approaches without large datasets.
Figure 1: We train geometry-informed neural networks to produce shapes satisfying geometric design requirements. For example, we generate parts that connect the cylindrical interfaces within the sketched design region depicted on the left. To highlight the user’s control over the problem and the solutions, we specify different additional requirements on the number of holes and surface smoothness. By complementing the design requirements with a diversity constraint, we can train a shape-generative model without data as illustrated in Figures 3 and 6.
@@ 1 Introduction
Fortunately, these disciplines are often equipped with formal problem descriptions, such as objectives and constraints. Previous works for PDEs [1], molecular science [2], and combinatorial optimization [3] demonstrate these can suffice to train models even in the absence of any data. The success of these data-free approaches motivates an analogous attempt in geometry, raising the question: Is it possible to train a shape-generative model on objectives and constraints alone, without relying on any data?
We address this question by introducing geometry-informed neural networks or GINNs. GINNs are trained to satisfy specified design constraints and to produce feasible shapes without any training samples. A GINN solves a topology optimization problem using neural fields, which offer detailed, smooth, and topologically flexible geometry representations, while being compact to store. This setup is analogous to physics-informed neural networks but with a high number of varied constraints: differential, integral, geometrical, and topological.
In contrast to both physics-informed neural networks and classical topology optimization, GINNs allow to generate multiple solutions by enforcing solution diversity as an explicit constraint. This is of high interest when applied to problems with solution multiplicity, e.g., induced by under-determinedness or near-optimality common in geometry problems. To connect back to our research question: with GINNs we can train neural fields that satisfy user-specified design constraints, and by adding diversity as an explicit constraint, we can generate a multiplicity of solutions. GINNs can thus be used as shape-generative models trained purely on constraints and without data.
We take several steps to demonstrate GINNs experimentally. We formulate a tractable learning problem using constrained optimization and by converting constraints into differentiable losses. After solving several introductory problems, we proceed to a realistic 3D engineering design problem. Figure 1 illustrates this task of designing a jet-engine lifting bracket, or geometrically – connecting cylindrical interfaces within the given design region. We show different GINNs trained with various additional smoothness and topology requirements. Figure 6 shows a GINN trained on the same task but with an additional diversity constraint. Surprisingly, this induces a structured latent space, with generalization capacity and interpretable directions.
We show that training shape-generative networks using constraints and objectives without data is a feasible learning strategy, paving the way for new generative design approaches without large datasets. Our main contributions are summarized as follows:
@@
    We apply GINNs to several validation problems and a realistic 3D engineering design problem, showing control over geometrical and topological properties. 111 Code available at https://github.com/ml-jku/GINNs-Geometry-informed-Neural-Networks
@@ 2.1 Theory-informed learning
Theory-informed learning has introduced a paradigm shift in scientific discovery by using scientific knowledge to remove physically inconsistent solutions and reducing the variance of a model [4]. Such knowledge can be included in the model via equations, logic rules, or human feedback [5, 6, 7]. Geometric deep learning [8] introduces a principled way to characterize problems based on symmetry and scale separation principles, e.g. group equivariances or physical conservation laws.
Notably, most works operate in the typical deep learning regime, i.e., with an abundance of data. However, in theory-informed learning, training on data can be replaced by training with objectives and constraints. More formally, one searches for a solution $f$ minimizing the objective $o(f)\text{ s.t. }f\in\mathcal{K}$, where $\mathcal{K}$ defines the feasible set in which the constraints are satisfied. For example, in Boltzmann generators [2], $f$ is a probability function parameterized by a neural network to approximate an intractable target distribution. Another example is combinatorial optimization where $f\in\{0,1\}^{N}$ is often sampled from a probabilistic neural network [9, 3, 10].
@@
(PINNs) [1] are a prominent example of neural optimization. In PINNs, $f$ is a function that must minimize the violation $o$ of a partial differential equation (PDE), the initial and boundary conditions, and, optionally, some measurement data. Since PINNs can incorporate noisy data and are mesh-free, they hold the potential to overcome the limitations of classical mesh-based solvers for high-dimensional, parametric, and inverse problems. This has motivated the study of the PINN architectures, losses, training, initialization, and sampling schemes [11]. We further refer to the survey by Karniadakis et al. [12].
Same as PINNs, GINNs use neural fields to represent the solution. Consequentially, we also observe that some best practices of training PINNs [11] transfer to training GINNs. However, PINNs may suffer from ill numerical properties due to minimizing the squared residual of the strong-form different to classical PDE solvers [13, 14]. In contrast, GINNs share the same underlying formulation and numerical properties as classical topology optimization methods. In addition to a high number of various constraints (differential, integral, geometrical, and topological), geometric problems often require solution multiplicity, motivating the generative extension.
@@ 2.2 Neural fields
A neural field (NF) (also coordinate-based NN, implicit neural representation (INR)) is a NN (typically a multilayer-perceptron) representing a function $f:x\mapsto y$ that maps a spatial and/or temporal coordinate $x$ to a quantity $y$. Compared to discrete representations, NFs are significantly more memory-efficient while providing higher fidelity, continuity, and access to automatic differentials. They have seen widespread success in representing and generating a variety of signals, including shapes [15, 16, 17], scenes [18], images [19], audio, video [20], and physical quantities [1]. For a more comprehensive overview, we refer to the survey by Xie et al. [21].
@@
(INSs) represent geometries through scalar fields, such as occupancy [17, 16] or signed-distance [15, 22]. In addition to the properties of NFs, INSs also enjoy topological flexibility supporting shape reconstruction and generation. We point out the difference between these two training regimes. In the generative setting, the training is supervised on the ground truth scalar field of every shape [15, 16, 17]. However, in surface reconstruction, i.e., finding a smooth surface from a set of points measured from a single shape, no ground truth is available and the problem is ill-defined  [22, 23].
@@
methods have been proposed to counter the ill-posedness in geometry problems. These include leveraging ground-truth normals [24] and curvatures [25], minimal surface property [24], and off-surface penalization [20]. A central effort is to achieve the distance field property of the scalar field for which many regularization terms have been proposed: eikonal loss [26], divergence loss [27], directional divergence loss [28], level-set alignment [29], or closest point energy [30]. The distance field property can be expressed as a PDE constraint called eikonal equation $|\nabla{f(x)}|=1$, establishing a relation of regularized INSs to PINNs [26].
@@
In addition to explicit loss terms, the architecture, initialization, and optimizer can also limit or bias the learned shapes. For example, typical INSs are limited to watertight surfaces without boundaries or self-intersections [31, 32]. ReLU networks are limited to piece-wise linear surfaces and together with gradient descent are biased toward low frequencies [33]. Fourier-feature encoding [33], sine activations [20], and wavelet activations [34] allow to control this frequency bias. Similarly, initialization techniques are important to converge toward desirable optima [20, 22, 27, 11].
@@
Generative modeling [35, 36, 37, 38] is almost exclusively performed in a data-driven (i.e., supervised) setting to capture and sampling from the underlying data-distribution. However, notable exceptions exist.
@@
[2] are a prominent example of data-free generative models. They are trained to capture the Boltzmann distribution associated with an energy landscape. In the generative setting, GINNs also learn a distribution minimizing an energy as an implicit combination of constraint violations and objectives. However, Boltzmann generators avoid mode-collapse using an entropy-regularizing term which presupposes invertibility making them not directly applicable to function spaces. Instead, GINNs use a more general diversity term to hinder mode-collapse over the function space of shapes.
@@
allow for generative modeling of functions. By conditioning a base network $F$ on a modulation (i.e, latent) variable $z$, a conditional NF encodes multiple fields simultaneously: $f(x)=F(x;z)$. The different choices of the conditioning mechanism lead to a zoo of architectures, including input concatenation [15], hypernetworks [39], modulation [40], and attention [41]. These can be classified into global and local mechanisms, which also establishes a connection between conditional NFs and operator learning [42]. For more detail we refer to Xie et al. [21], Rebain et al. [41], Wang et al. [42].
@@
refers to computational methods that automatically conduct design exploration under constraints set by designers [43]. It holds the potential of streamlining innovative solutions, e.g., in material design, architecture, or engineering. In particular, GINNs can be seen as solving the general task of topology optimization – finding the material distribution that minimizes a specified objective subject to constraints. However, while classical methods optimize a single shape directly, we optimize a GINN that generates diverse feasible shapes. This encourages design space exploration and supports downstream tasks, while allowing to incorporate even sparse data samples, if available. While generative design datasets are not abundant, deep learning has previously shown promise in material design and topology optimization. For more detail, we refer to surveys on generative models in engineering design [44] and topology optimization via machine learning [45].
@@ 3 Method
Consider the metric space $(d,\mathcal{F})$ of functions, such as those representing a shape or a PDE solution. Let the set of constraints define the feasible set $\mathcal{K}=\{f\in\mathcal{F}|c_{i}(f)=0,i=1..m\}$. Additionally, let the geometric problem be equipped with an objective $o:\mathcal{F}\mapsto\mathbb{R}$. Selecting the objectives and constraints of a geometric nature lays the foundation for a GINN, which is trained to produce an optimal feasible solution by solving $\min_{f\in\mathcal{K}}o(f)$. A key feature of geometric problems is that one is often interested in finding different near-optimal solutions, for example, due to incompleteness, uncertainty, or under-determinedness in the problem specification (e.g. see Figure A.4). This motivates making GINN generate a set of sufficiently diverse near-optimal solutions $S\subset\mathcal{K}$:
$\min_{\begin{subarray}{c}S\subset\mathcal{K}\\ \delta(S)\geq\delta_{\text{min}}\end{subarray}}O(S)\ .$ (1)
$O(S)$ aggregates the objectives $o(f)$ of all solutions $f\in S$ and $\delta$ captures some intuitive notion of diversity. It is yet another constraint, however it acts on the entire solution set instead of each solution separately. Section 3.1 first discusses representing shapes as functions, in particular, neural fields, and formulating differentiable constraints. In Section 3.2, we generalize to representing and finding diverse solutions using conditional neural fields.
@@ 3.1 Finding a solution
Prescribed normal $n(x)=\bar{n}(x)\ \forall x\in\mathcal{I}$ $\frac{\nabla f(x)}{||\nabla f(x)||}=\bar{n}(x)\ \forall x\in\mathcal{I}$ $\int_{\mathcal{I}}\left[\frac{\nabla f(x)}{||\nabla f(x)||}-\bar{n}(x)\right]^{2}\operatorname{d}\!{x}$
@@
Table 1: Geometric constraints used in our main experiment. The shape $\Omega$ and its boundary $\partial\Omega$ are represented implicitly by the (sub-)level set of the function $f$. The shape must be contained within the design region $\mathcal{E}\subseteq\mathcal{X}$ and attach to the interface $\mathcal{I}\subset\mathcal{E}$ with a prescribed normal $\bar{n}(x)$. Other interesting constraints are listed in Table 5.
@@
Let $f:\mathcal{X}\mapsto\mathbb{R}$ be a continuous scalar function on the domain $\mathcal{X}\subset\mathbb{R}^{n}$. The sign of $f$ implicitly defines the shape $\Omega=\left\{x\in\mathcal{X}|f(x)\leqslant 0\right\}$ and its boundary $\partial\Omega=\left\{x\in\mathcal{X}|f(x)=0\right\}$. We use a NN $f=f_{\theta}$ with parameters $\theta$ to represent the implicit function, i.e. an implicit neural shape, due to its memory efficiency, continuity, and differentiability. Nonetheless, the GINN paradigm extends to other representations, as demonstrated in Section 4.2.
@@
To perform gradient-based optimization, we must first ensure each constraint can be written as a differentiable constraint violation $c_{i}:\mathcal{F}\mapsto\mathbb{R}$. A geometric constraint has the general form $c_{i}(\Omega,\partial\Omega)=0$. By representing the shape and its boundary as the (sub-)level-set of the function $f$, the constraints on the sets can be translated into constraints on $f$. This in turn allows to formulate differentiable constraint violations $c_{i}$, although this choice is not unique. Table 1 shows several examples using the constraints from our main experiment. Some losses are straightforward and some have been previously demonstrated as regularization terms for INSs (see Section 2.2). Section 4.1 discusses two complex losses in more detail: connectedness and smoothness.

### v4

@@ Abstract
Geometry is a ubiquitous tool in computer graphics, design, and engineering. However, the lack of large shape datasets limits the application of state-of-the-art supervised learning methods and motivates the exploration of alternative learning strategies. To this end, we introduce geometry-informed neural networks (GINNs) – a framework for training shape-generative neural fields without data by leveraging user-specified design requirements in the form of objectives and constraints. By adding diversity as an explicit constraint, GINNs avoid mode-collapse and can generate multiple diverse solutions, often required in geometry tasks. Experimentally, we apply GINNs to several problems spanning physics, geometry, and engineering design, showing control over geometrical and topological properties, such as surface smoothness or the number of holes. These results demonstrate the potential of training shape-generative models without data, paving the way for new generative design approaches without large datasets.
geometry, implicit neural representation, neural fields, theory-informed learning, geometric deep learning, physics-informed neural networks, generative design
Figure 1: We train geometry-informed neural networks to produce shapes satisfying geometric design requirements. For example, we generate parts that connect the cylindrical interfaces within the sketched design region depicted on the left. To highlight the user’s control over the problem and the solutions, we specify different additional requirements on the number of holes and surface smoothness. By complementing the design requirements with a diversity constraint, we can train a shape-generative model without data as further illustrated in Figures 3, 5 and 6.
@@ 1 Introduction
Fortunately, these disciplines are often equipped with formal problem descriptions, such as objectives and constraints. Previous works for PDEs (Raissi et al., 2019), molecular science (Noé et al., 2019), and combinatorial optimization (Bengio et al., 2021) demonstrate these can suffice to train models even in the absence of any data. The success of these data-free approaches motivates an analogous attempt in geometry, raising the question: Is it possible to train a shape-generative model on objectives and constraints alone, without relying on any data?
We address this question by introducing geometry-informed neural networks or GINNs. GINNs are trained to satisfy specified design requirements and to produce feasible shapes without any training samples. A GINN solves a structural optimization problem using neural fields, which offer detailed, smooth, and topologically flexible geometry representations, while being compact to store. This setup is analogous to physics-informed neural networks but with a high number of varied constraints: differential, integral, geometrical, and topological.
In contrast to both physics-informed neural networks and classical structural optimization, GINNs allow to generate multiple solutions by enforcing solution diversity as an explicit constraint. This is of high interest when applied to problems with solution multiplicity, e.g., induced by under-determinedness or near-optimality common in geometry problems. Connecting back to our research question, the proposed framework allows us to train neural fields that satisfy user-specified design constraints, and by adding diversity as an explicit constraint, we can generate a multiplicity of solutions. GINNs can thus be used as shape-generative models trained purely on constraints and without data.
We take several steps to demonstrate GINNs experimentally. We formulate a tractable learning problem using constrained optimization and by converting constraints into differentiable losses. We demonstrate the generality of the framework by solving several tasks, including an under-determined PDE, geometric optimization, and engineering design (Figures 4, 5). Figure 1 illustrates the task of designing a jet-engine lifting bracket, or geometrically – connecting cylindrical interfaces within the given design region. We show different GINNs trained with various smoothness and topology requirements, illustrating the robustness of the constrained optimization approach and the user’s control over the framework. Figure 6 shows a GINN trained on the same task but with an additional diversity constraint. Surprisingly, this constraint induces a structured latent space, with generalization capacity and interpretable directions.
We show that training shape-generative networks using constraints and objectives without data is a feasible learning strategy, paving the way for new generative design approaches without large datasets. Our main contributions are summarized as follows [^1] Code is available at https://github.com/ml-jku/GINNs-Geometry-informed-Neural-Networks:
@@
    We apply GINNs to several problems spanning physics, geometric optimization, and engineering design, investigating the user’s control, the robustness of the method, key framework choices, and emerging latent space structure.
@@ 2.1 Theory-Informed Learning
Theory-informed learning has introduced a paradigm shift in scientific discovery by using scientific knowledge to remove physically inconsistent solutions and reducing the variance of a model (Karpatne et al., 2017). Such knowledge can be included in the model via equations, logic rules, or human feedback (Dash et al., 2022; Muralidhar et al., 2018; Von Rueden et al., 2021). Geometric deep learning (Bronstein et al., 2021) introduces a principled way to characterize problems based on symmetry and scale separation principles, e.g. group equivariances or physical conservation laws.
Notably, most works operate in the typical deep learning regime, i.e., with an abundance of data. However, in theory-informed learning, training on data can be replaced by training with objectives and constraints. More formally, one searches for a solution $f$ minimizing the objective $o(f)\text{ s.t. }f\in\mathcal{K}$, where $\mathcal{K}$ defines the feasible set in which the constraints are satisfied. For example, in Boltzmann generators (Noé et al., 2019), $f$ is a probability function parameterized by a neural network to approximate an intractable target distribution. Another example is combinatorial optimization where $f\in\{0,1\}^{N}$ is often sampled from a probabilistic neural network (Bello et al., 2016; Bengio et al., 2021; Sanokowski et al., 2024).
@@
(PINNs) (Raissi et al., 2019) are a prominent example of neural optimization. In PINNs, $f$ is a function that must minimize the violation $o$ of a partial differential equation (PDE), the initial and boundary conditions, and, optionally, some measurement data. Since PINNs can incorporate noisy data and are mesh-free, they hold the potential to overcome the limitations of classical mesh-based solvers for high-dimensional, parametric, and inverse problems. This has motivated the study of the PINN architectures, losses, training, initialization, and sampling schemes (Wang et al., 2023). We further refer to the survey by Karniadakis et al. (2021).
Same as PINNs, GINNs use neural fields to represent the solution. Consequentially, we also observe that some best practices of training PINNs (Wang et al., 2023) transfer to training GINNs. However, PINNs may suffer from ill numerical properties due to minimizing the squared residual of the strong-form different to classical PDE solvers (Rathore et al., 2024; Ryck et al., 2024). In contrast, GINNs share the same underlying formulation and numerical properties as classical topology optimization methods. In addition to a high number of various constraints (differential, integral, geometrical, and topological), geometric problems often require solution multiplicity, motivating the generative extension.
@@ 2.2 Neural Fields
A neural field (NF) (also coordinate-based NN, implicit neural representation (INR)) is a NN (typically a multilayer-perceptron) representing a function $f:x\mapsto y$ that maps a spatial and/or temporal coordinate $x$ to a quantity $y$. Compared to discrete representations, NFs are significantly more memory-efficient while providing higher fidelity, continuity, and access to automatic differentials. They have seen widespread success in representing and generating a variety of signals, including shapes (Park et al., 2019; Chen & Zhang, 2019; Mescheder et al., 2019), scenes (Mildenhall et al., 2021), images (Karras et al., 2021), audio, video (Sitzmann et al., 2020), and physical quantities (Raissi et al., 2019). For a more comprehensive overview, we refer to the survey by Xie et al. (2022).
@@
(INSs) represent geometries through scalar fields, such as occupancy (Mescheder et al., 2019; Chen & Zhang, 2019) or signed-distance (Park et al., 2019; Atzmon & Lipman, 2020). In addition to the properties of NFs, INSs also enjoy topological flexibility supporting shape reconstruction and generation. We point out the difference between these two training regimes. In the generative setting, the training is supervised on the ground truth scalar field of every shape. However, in surface reconstruction, i.e., finding a smooth surface from a set of points measured from a single shape, no ground truth is available and the problem is ill-defined  (Atzmon & Lipman, 2020; Berger et al., 2016).
@@
methods have been proposed to counter the ill-posedness in geometry problems. These include leveraging ground-truth normals (Atzmon & Lipman, 2021) and curvatures (Novello et al., 2022), minimal surface property (Atzmon & Lipman, 2021), and off-surface penalization (Sitzmann et al., 2020). A central effort is to achieve the distance field property of the scalar field for which many regularization terms have been proposed: eikonal loss (Gropp et al., 2020), divergence loss (Ben-Shabat et al., 2022), directional divergence loss (Yang et al., 2023), level-set alignment (Ma et al., 2023), or closest point energy (Marschner et al., 2023). The distance field property can be expressed as a PDE constraint called eikonal equation $|\nabla{f(x)}|=1$, establishing a relation of regularized INSs to PINNs (Gropp et al., 2020).
@@
In addition to explicit loss terms, the architecture, initialization, and optimizer can also limit or bias the learned shapes. For example, typical INSs are limited to watertight surfaces without boundaries or self-intersections (Chibane et al., 2020; Palmer et al., 2022). ReLU networks are limited to piece-wise linear surfaces with sharp vertices and edges and, together with gradient descent, are biased toward low frequencies (Tancik et al., 2020). Fourier-feature encoding (Tancik et al., 2020), sine activations (Sitzmann et al., 2020), and wavelet activations (Saragadam et al., 2023) allow for controlling this frequency bias. Similarly, initialization techniques are important to converge toward desirable optima (Sitzmann et al., 2020; Atzmon & Lipman, 2020; Ben-Shabat et al., 2022; Wang et al., 2023).
@@
Generative modeling (Kingma & Welling, 2013; Goodfellow et al., 2014; Rezende & Mohamed, 2015; Tomczak, 2021) is almost exclusively performed in a data-driven (i.e., supervised) setting to capture and sample from the underlying data distribution. However, notable exceptions exist.
@@
(Noé et al., 2019) are a prominent example of data-free generative models. They are trained to capture the Boltzmann distribution associated with an energy landscape. In the generative setting, GINNs also learn a distribution minimizing an energy as an implicit combination of constraint violations and objectives. However, Boltzmann generators avoid mode-collapse using an entropy-regularizing term, which presupposes invertibility, making them not directly applicable to function spaces. Instead, GINNs use a more general diversity term to hinder mode-collapse over the function space of shapes.
@@
allow for generative modeling of functions. By conditioning a base network $F$ on a modulation (i.e., latent) variable $z$, a conditional NF encodes multiple fields simultaneously: $f(x)=F(x;z)$. The different choices of the conditioning mechanism lead to a zoo of architectures, including input concatenation (Park et al., 2019), hypernetworks (Ha et al., 2017), modulation (Mehta et al., 2021), and attention (Rebain et al., 2022). These can be classified into global and local mechanisms, which also establishes a connection between conditional NFs and operator learning (Wang et al., 2024). For more detail, we refer to Xie et al. (2022); Rebain et al. (2022).
@@
refers to computational methods that automatically conduct design exploration under constraints set by designers (Jang et al., 2022). It holds the potential of streamlining innovative solutions, e.g., in material design, architecture, or engineering. GINNs can be seen as solving the general task of topology optimization – finding the material distribution that minimizes a specified objective subject to constraints. However, while classical methods optimize a single shape directly, we optimize a model that generates diverse feasible shapes. This supports design space exploration and downstream tasks while allowing to incorporate sparse data samples, if available. While generative design datasets are not abundant, deep learning has shown promise in material design and topology optimization. For more detail, we refer to surveys on generative models in engineering design (Regenwetter et al., 2022) and topology optimization via machine learning (Shin et al., 2023).
@@ 3 Method
Consider the metric space $(\mathcal{F},d)$ of functions, such as those representing a shape or a PDE solution. Let the set of constraints define the feasible set $\mathcal{K}=\{f\in\mathcal{F}|c_{i}(f)=0,i=1..m\}$. Additionally, let the problem be equipped with an objective $o:\mathcal{F}\mapsto\mathbb{R}$. Selecting the objectives and constraints of a geometric nature lays the foundation for a GINN, which is trained to produce an optimal feasible solution by solving $\min_{f\in\mathcal{K}}o(f)$. A key feature of geometric problems is that one is often interested in finding different near-optimal solutions, for example, due to incompleteness, uncertainty, or under-determinedness in the problem specification (e.g., see Figure B.4). This motivates making GINN generate a set of sufficiently diverse near-optimal solutions $S\subset\mathcal{K}$:
$\min_{\begin{subarray}{c}S\subset\mathcal{K}\\ \delta(S)\geq\bar{\delta}\end{subarray}}O(S)\ .$ (1)
$O(S)$ aggregates the objectives $o(f)$ of all solutions $f\in S$ and $\delta$ captures some intuitive notion of diversity. While it can be treated as another constraint, it acts on the entire solution set instead of each solution separately. Section 3.1 first discusses representing shapes as functions, in particular, neural fields, and formulating differentiable constraints. In Section 3.2, we generalize to representing and finding diverse solutions using conditional neural fields.
@@ 3.1 Finding a Solution
Prescribed normal $n(x)=\bar{n}(x)\ \forall x\in\mathcal{I}$ $\nabla f(x)=\bar{n}(x)\ \forall x\in\mathcal{I}$ $\int_{\mathcal{I}}\left[\nabla f(x)-\bar{n}(x)\right]^{2}\operatorname{d}\!{x}$
@@
Table 1: Geometric constraints used in several experiments. The shape $\Omega$ and its boundary $\partial\Omega$ are represented implicitly by the (sub-)level set of the function $f$. The shape must be contained within the design region $\mathcal{E}\subseteq\mathcal{X}$ and attach to the interface $\mathcal{I}\subset\mathcal{E}$ with a prescribed unit normal $\bar{n}(x)$. Other interesting constraints are listed in Table 8.
@@
Let $f:\mathcal{X}\mapsto\mathbb{R}$ be a continuous scalar function on the domain $\mathcal{X}\subset\mathbb{R}^{n}$. The sign of $f$ implicitly defines the shape $\Omega=\left\{x\in\mathcal{X}|f(x)\leq 0\right\}$ and its boundary $\partial\Omega=\left\{x\in\mathcal{X}|f(x)=0\right\}$. We use a NN $f=f_{\theta}$ with parameters $\theta$ to represent the implicit function, i.e., an implicit neural shape, due to its memory efficiency, continuity, and differentiability. Nonetheless, the GINNs extend to other representations, as shown in Section 4.2.
@@
To perform gradient-based optimization, we must first ensure each constraint can be written as a differentiable constraint violation $c_{i}:\mathcal{F}\mapsto\mathbb{R}$. A geometric constraint has the general form $c_{i}(\Omega,\partial\Omega)=0$. By representing the shape and the boundary implicitly via the function $f$, the constraints on the sets can be translated into constraints on $f$. This in turn allows to formulate differentiable constraint violations $c_{i}$, although this choice is not unique. Table 1 shows several examples using the constraints from our experiments. Some losses are straightforward, and some have been previously demonstrated as regularization terms for INSs (see Section 2.2). Section 4.1 discusses two complex losses in more detail: connectedness and smoothness.